does the ring Z2[x]/<(x^8)+1> have nilpotent elements? justify.
1
Expert's answer
2020-10-12T18:09:49-0400

An element "x" of a ring "R" is called nilpotent if there exists some positive integer "n" such that "x^n=0".


Polynomial remainder theorem implies that each element of "\mathbb Z_2[x]/\langle x^8+1\rangle" is of the form "[f(x)]=f(x)+\langle x^8+1\rangle=f(x)+(x^8+1)\mathbb Z_2[x]", where "\deg f(x)<\deg(x^8+1)=8". Thus, "[0]=\langle x^8+1\rangle=(x^8+1)\mathbb Z_2[x]." Since "1+1=0" in the ring "\mathbb Z_2[x]" and "x^8+1\in\langle x^8+1\rangle=[0]", "[(x^4+1)]^2=[(x^4+1)^2]=[x^8+x^4+x^4+1]=[x^8+(1+1)x^4+1]=[x^8+1]=[0]."

Consequently, "[x^4+1]" is a nilpotent element.